Prev Next

SW&IT Meeting Minutes 2020-09-28

Attendees: Carl, David, Igor, Matthieu, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200921
LR20200921 - Bec, Matthieu 40.0
LR20200921 - Holmberg, Carl 40.0
LR20200921 - Edmeades, David 40.0
LR20200921 - Sola, Igor 0.0
LR20200921 - Hooper, Stephen 80.0
LR20200921 - Bryant, Trey 0.0
LR20200921 - Zhang, Yang 40.0
LR20200921 - Zhang, Xianyu 40.0

News from management

  • 1-3 years budget plans, need input from IT
  • MOEBIUS - and plans to use IRTC camera for Luci

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • lbc / robs - effort continues

Team reports

MB

  • Last week
    • OC, Management
    • ATT mostly
    • ROWG document
    • review different ITs (vx socket err, PMCGUI, NUC satellite, robs du, etc.)
  • This week
    • ROWG meeting
    • leadership workshop Tue/Wed 8-12am
    • ATT mostly

CH

  • Last week:
    • AGW: JPH binary encoder integration.
      • Provided him with my PMAC firmware notes on grey encoder interface.
      • He has successfully tested the encoder/interface board using the Pwen PC software.
      • We will likely need to at least edit the current PMAC .cfg file to reflect the setting for the new encoder and word size.
    • DIMM: Cloned the repo, reviewed the existing wiki docs.
    • DMS: Backfilled Mongodb with previously undelivered iMonnit data. Set sensor sample rates from 2 hours back to 10 minutes.
    • LBC: 64 bit transition.
      • Attended zoom session to review functionality of LBC tools with OK, YZ, SH.
      • Prototyped approaches to adding timestamp label back into DS9 from RB_Science_cmh_dev.py via XPA, but probably moot.
      • Retested and patched fits verification function in RB_Science_cmh_dev.py.
      • Still suffering seemingly random crashes with both old and new script versions. Running in python trace mode to see if it will illuminate the source of the errors.
      • Testing with revolving set of 10 recent LBC fits files fails to force an RB_Science_cmh_dev.py error on robs user file system.
      • Created script to load test data. Cleared out excess test FITS files from robs which were robbing the system of local file space.
    • LBTI: AHS client fix for cryopump command status deployed.

  • This week:
    • AGW: On Monday JPH will re-review my notes on existing PMAC firmware to determine if we need to change the scaling to convert encoder interface board output to 32 bit Signed format. JPH will test Serial SSI data Gray-to-binary convert control word to see if we need to unset the gray-to-binary-code flag. From these results we can look at what code patches are needed, and rebuild the firmware. I don’t yet know enough to say if that will happen this week.
    • DIMM: Follow SH progress, attempt build in parallel.
    • DMS: iMonnit sensor settings check. Screen captures of various sensor type settings pages sent to Peter Gray for review. Will correct settings if needed.
    • LBC 64 bit transition
      • Retry RB_Science.py revolving fits test using NFS directory.
      • Participate in any follow up discussions.

DE

  • Zoom security updates
    • Mostly smooth, hiccups based on user behavior re: Zoom login
    • H.323 bridge solid
  • Veeam ticket opened
    • They think we have a bug, and have supplied a patched DLL
  • Synology remote access granted to mountain
  • Moved Vault VM to VMHost1
    • Last VM to move
    • VMHost2 is empty
    • Hooray, no more Flash
  • Office visit to swap disk, get PEPSI server on its way to the summit
Next:
  • XZ volume on mountan for AO logs
  • Veeam DLL fix (should not involve downtime)
  • Cisco
  • Probably more handholding about Zoom
    • Gino is making up wild theories, but has not responded to my request for a screenshot or the instructions that UITS sent
    • Haven't heard anything from Robert so I'm sure he has no idea what's going on

XZ

  • Last week
    • Build a TCS and MCSPU simulator on my laptop, thanks to Igor for helping
    • AdSec TSS with high wind analyze
    • ATT SW list
    • New "bump" on sxadsec
    • AO support
  • This week
    • Handle shell rested but dome open in AOS?
    • ATT SW and AdSec
    • Add new info in AdSec event script
    • Safety training
    • AO support

YZ:

  • Previous:
    • TMS test on Monday:
      • More problems, reflects more unknown of the new TMS inverse kinematics model and perhaps more misunderstandings.
      • New approach (and patches) proposed, need to study them and implement them. However without a solid theoretical understanding and certain mathematical proof, it is hard to believe this will be the final correct version.
    • LBC 64-bit operations test and plan.
    • AO-Cravat (meeting and plan).
  • Next:
    • TMS work as usual.
    • AO-Cravat work as usual.
    • LBC 64-bit operations to continue (several remaining issues, some small some big, e.g., TCSSendWavefront has ICE warnings, RB_Science does not work very well).
    • Several other misc activities if time permits (issue trackes, LBC/MODS misc works).

SH

  • DIMM (Tue, Sat, Sun)
    • Installed on Saturday, removed on Sunday
    • Software seems to work Ok, hardware (network now) was problematic
    • LeRoy moved the disk and EDT card to the eBay pruchase. I configured network, and run network tests to make sure it is good.
    • Josh had some trouble with software
  • VxWorks tracing (Fri/Sat)
  • VPN LDAP (Fri,Wed)
  • 64-bit LBC meeting (Thu)
  • ROWG and UPS sizing (Thu)
  • Wired network (Fri,Tue)
  • Siemens PLC licensing (Wed, Tue, Mon)
  • rm507 issues (Tue, Wed, Thu)

next:

  • DIMM testing
  • LeRoy PEPSI servers
  • Github (webapps, etc.)
  • PMC network (checks)
  • LDAP VPN

IS

Last week:
  • Safety Training
  • PIT
    • Last two days of the week dedicated to the Close for Proposals opening and closing dates, and listing of current proposals per partner to the pit admin tool. Still some more work to do.
  • LUCI
    • Along with Xianyu we built a TCS test environment. I further configure mine to work with my LUCI test environment. Both VMs can run in my laptop.
    • Test with Elliot lab webio but decided to wait for next time one of the LUCI's is back in the lab as we need to access the WebIO behind the electronics rack.

This week:

  • Safety Training
  • PIT
    • the Close for Proposals opening and closing dates, and listing of current proposals per partner to the pit admin tool
  • LUCI
    • Try to do the same I did for the TCS VM in a docker container, and finish the to write the doc.
    • Test with Elliot lab webio but decided to wait for next time one of the LUCI's is back in the lab as we need to access the WebIO behind the electronics rack.
    • Add more debugging information in Journalizer DB storing logic. Trying to figure out why sometimes the mask config file doesn't upload.
Topic revision: r13 - 28 Sep 2020, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback